Wilma drove and average speed of 60mph from her hom in city A to visit her sister in city B. She stayed in City b 15 hours, and on the trip back averaged 40mph. She returned home 49 hours after leaving. Whave have to determine how many miles x is it form City A to City B. In terms of x, express wilma's driving time from A to B. In terms o fher, her return driving time is? Her total driving time is? (trip time minus visiting time)
Answer by checkley71(8403)

60T+40[(49-15)-T]
60T=40*34-40T
100T=1360
T=13.6 HOURS FOR THE TRIP FROM A TO B.
34-13.60=20.4 H0URS FOR THE RETURN TRIP FROM B TO A.
PROOF
60*13.6=40*20.4
816=816 MILES BETWEEN A & B.
13.6+20.4=34 HOURS IS HER TOTAL DRIVING TIME.